A 20 g sample of ice at -10 °C is mixed with 100 g water at 80 °C. Calculate the final temperature of the mixture.

1
Expert's answer
2021-11-02T02:24:30-0400

Let the final temperature be T

Heat released by water = mass "\times" specific heat capacity "\times" temperature change

The specific heat of liquid water is 4.18 J/g°C

"= 100 \times 4.18 \times (80 - T)"

Heat absorbed by ice on melting = mass "\times" heat of fusion

The heat of fusion of ice is 335 J/g

"= 20 \times 335 = 6700 \;J"

Heat absorbed by melted ice = mass "\times" specific heat capacity "\times" temperature change

"= 20 \times 4.18 \times (T - (-10)) = 83.6(T+10)"

Total heat released = total heat absorbed

"100 \times 4.18 \times (80 - T) = 6700 + 83.6(T+10) \\

33440 -418T = 6700 +83.6T + 836 \\

25904 = 501.6T \\

T = \frac{25904}{501.6} \\

T = 51.64 \; °C"

Answer: 51.64 °C